OpenStudy (anonymous):

Carlos solved the system of equations below. What mistake did he make in his work? 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x + 10 + 4x = 10 5x + 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5 He should have substituted 5 + 2x He did not distribute correctly He combined like terms incorrectly, it should have been 4x instead of 5x He made no mistake

OpenStudy (anonymous):

The mistake starts at the part when -2(5-2x) The consequent equation should be -10+4x and not +10+4x

OpenStudy (anonymous):

so he should have substituted 5 +2x?

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

No. y = 5 − 2x is correct.

OpenStudy (anonymous):

so He did not distribute correctly is it right

OpenStudy (anonymous):

He distributed correctly. Take a look at this. 2x + y = 5 2x + y -2x = 5 -2x y = 5 - 2x

OpenStudy (anonymous):

cause its not the 3rd choice because i chose that last time and it was wrong

OpenStudy (anonymous):

so he made no mistake then

OpenStudy (anonymous):

He did make a mistake but it's not in all four of the choices.

OpenStudy (anonymous):

I've already written the mistake in my first reply here.

OpenStudy (anonymous):

i know but thats not in the choices of the answer options

OpenStudy (anonymous):

Since there's a mistake (4th choice out), and the 1st and 3rd choices aren't correct too, so the 2nd choice should be the best answer.

OpenStudy (anonymous):

okay so he didnt distribute correctly

OpenStudy (anonymous):

I do not know what "distribute" here means anyway. Seems to me it isn't like converting the wrong sign.

OpenStudy (anonymous):

soo.. it would be the second one?

OpenStudy (goformit100):

Yeah

OpenStudy (anonymous):

Yes. The only logical answer after deducting all the confirmed-to-be wrong ones.